A certain amount was to be distributed among P,Q & R in the ratio of 12:13:15 but was erroneously distributed in the ratio of 2:3:5. As a result, R got Rs. 585 more. What was the amount distributed ?
Answer by mananth(16946)

let the amount distributed be x
1 part = 1/40
So R gets 15x/40
=3x/8
Actual distribution
each part = 1/10 = 4680
R got x/2
x/2 - 3x/8 =585
x/8 = 585
x= 8*585 = 4680